# Implementation of Time-Varying Controllers for a Nonholonomic Mobile Robot: Experimental Studies

### v1

@@ 2 Stabilizing Controls for the Unicycle Kinematics
In this paper, we consider the well-known kinematic model of a unicycle rolling on the $(x_{1},x_{2})$-plane, described by the following system of differential equations (see, e.g., [4]):
@@
where the vector fields $f_{1},f_{2}:{\mathbb{R}}^{3}\to{\mathbb{R}}^{3}$ are defined as $f_{1}(x)=\big(\cos x_{3},\sin x_{3},0\big)^{\top}$, $f_{2}(x)=\big(0,0,1\big)^{\top}$. Here, $(x_{1},x_{2})$ are the coordinates of the contact point of the wheel, and $x_{3}$ is the angle between the plane of the wheel and the $x_{1}$-axis. The controls $u_{1}$ and $u_{2}$ describe the translational and angular velocities of the unicycle, respectively. To simplify the presentation, all variables of this control system are assumed to be dimensionless. For the purpose of control design, we introduce the matrix ${\cal F}(x)$ composed of the vector fields $f_{1}(x)$, $f_{2}(x)$ and their Lie bracket $f_{1},f_{2}](x)=\frac{\partial f_{2}(x)}{\partial x}f_{1}(x)-\frac{\partial f_{1}(x)}{\partial x}f_{2}(x)$: F(x) = (f_1(x),f_2(x),[f_1,f_2](x)) = (cosx30 sinx3sinx30 - cosx30 1 0 ). This matrix is nonsingular for all $x\in{\mathbb{R}}^{3}$, and its inverse is given by
${\cal F}^{-1}(x)=\begin{pmatrix}\cos x_{3}&\sin x_{3}&0\\ 0&0&1\\ \sin x_{3}&-\cos x_{3}&0\end{pmatrix}$. (2)
The nonsingularity of ${\cal F}(x)$ implies that the driftless control-affine system (1) satisfies the Lie Algebra Rank Condition (LARC): span (f_1(x),f_2(x),[f_1,f_2](x)) = R^3 for all   x∈R^3. Hence, by the Chow–Rashevsky theorem, system (1) is completely controllable in ${\mathbb{R}}^{3}$. For a given real-valued potential function $V\in C^{2}({\mathbb{R}}^{3})$, the approximate motion of system (1) in the direction of the negative gradient of $V(x)$ is generated by control functions $u^{\omega}(x,t)=(u_{1}^{\omega}(x,t),u_{2}^{\omega}(x,t))^{\top}$ of the following form [8]:
@@ 3 Justification for the Lyapunov Function Candidate
A natural way of defining an optimal Lyapunov function candidate for system (1) in $X$ can be as follows: find a function $V\in C^{2}(X;{\mathbb{R}})$ that minimizes the cost (9) under the constraint that $V$ is positive definite in $X$. This optimization problem is highly nonlinear and nonconvex, and it poses significant analytical and computational challenges. In particular, the set of positive definite functions is nonconvex in any reasonable function space, and the functional (9) exhibits singular behavior at critical points of $V$ due to the presence of $\nabla V$. Note that, in the case $U={\mathbb{R}}^{2}$, the functional $J_{X}[V]$ is invariant under multiplication of $V$ by a nonzero constant. Therefore, a solution to the above optimization problem (if it exists) is not unique without additional constraints. A comprehensive analysis of this general problem is beyond the scope of the present paper. Instead, we restrict our attention to estimating $J_{X}[V]$ within some positive definite quadratic forms.
@@
where $\alpha$ is a parameter. In particular, computing the cost (9) for $V_{\alpha}$ with $\alpha=2,4,10$, we obtain: J_X[V_2]≈0.1403,  J_X[V_4]≈0.0962,  J_X[V_10]≈0.0906. The obtained sequence of decreasing cost values justifies the use of the Lyapunov function candidates (13) for further experiments. It should be emphasized that quadratic forms cannot serve as control-Lyapunov functions for systems of the form (1) due to Brockett’s necessary stabilizability condition. Indeed, Brockett’s condition implies that the considered class of nonholonomic systems does not admit any continuously differentiable control-Lyapunov function.
@@ 4.1 Experimental Setup
Robotic Platform. The experiments were conducted using a TurtleBot3 Burger (TB3) – an affordable, differential-drive mobile robot (see Figure 1, left) [1]. The robot’s control and communication were handled by its onboard Raspberry Pi 4B, running Ubuntu Server 22.04.5 LTS (64-bit) and ROS 2 Humble. Mechanically, the TB3 features two off-centered, independently actuated wheels and a caster ball, thus the robot is not perfectly unicycle. However, a differential-drive robot can be modeled as unicycle by mapping the velocities of the left and right wheel $v_{l}$ and $v_{r}$ (linear velocities at the wheel edge) to the unicycle velocities $u_{1}$ and $u_{2}$ mentioned in (1). Let $d$ be the distance between the two wheels of the differential drive robot, and assume the robot’s center of rotation is at the midpoint between the wheels. Then, the unicycle’s translational and angular velocities, respectively $u_{1}$ and $u_{2}$, are given by $u_{1}=\frac{v_{r}+v_{l}}{2},\quad u_{2}=\frac{v_{r}-v_{l}}{d}.$ The inverse of this transformation is handled automatically by the diff_drive_controller of ROS 2. It accepts the unicycle-style Twist commands on /cmd_vel and computes the corresponding wheel velocities. The robot’s motion is constrained by its kinematics with the maximum translational velocity of 0.22 m/s and the maximum angular velocity of 2.84 rad/s. No onboard sensors were used for state estimation or control, but four reflective markers were mounted for the motion capture system.
Motion Capture System. Robot pose information was obtained using an OptiTrack motion capture system, consisting of eight infrared cameras and the Motive tracking software. As previously stated, four reflective markers were attached to the top of the TB3, allowing the estimation of its x- and y-position as well as the yaw orientation. The Motive software typically computes the center point of these markers and uses this as a reference pose. However, this point did not align precisely with the robot’s pivot point. To ensure accurate control, the reference pose was adjusted manually to match the true center of rotation (see Figure 1, right). We used a dedicated wireless access point to connect to the OptiTrack system via LAN, ensuring the TB3 and a remote laptop used for robot control could communicate in a known low-latency wireless network.
@@
Software Architecture. The control logic for both simulation and real-world experiments was implemented as a single ROS 2 node using the rclpy Python library. In real-world experiments, the controller subscribes to the robot pose data provided by the OptiTrack system via the NatNet protocol, whereas in simulations it subscribes to the standard /odom ROS 2 topic published by Gazebo. Control commands are published on the standard /cmd_vel topic, recognized by both the real-world TB3 and the Gazebo simulation. This allows us to use one and the same controller logic in different domains. Control commands are recomputed and sent as fast as possible which turned out to be every 0.5 milliseconds. Experimental data and controller parameters were recorded automatically for each trial. Plots were generated using Python’s matplotlib library, and all data were exported in .csv format for further analysis.
Simulation Environment. Simulation experiments were conducted using the predefined TB3 Gazebo environment provided in the Robotis e-Manual [2]. An empty world was used to replicate a simple, obstacle-free arena. The same TB3 model (Burger) as in real world was used to ensure consistency in dimensions, kinematics and dynamics. The model incorporates basic rigid body physics including link masses and joint properties allowing the robot to respond realistically to velocity commands although small inaccuracies in terms of friction and inertia compared to the real-world robot may exist. Due to the software architecture the same Python controller could be used.
@@ 4.2 Experimental Results
For all experiments, the initial state is set to $x^{0}=(-0.5,-0.5,0)^{\top}$, and the target state is $x^{*}=(0,0,0)^{\top}$. The robot stops automatically if it is in the vicinity of that goal point to obtain comparable convergence time results. The translational and angular velocities are governed by the control law (3), (5) with the frequency parameter $\omega=2\pi$ (i.e., $\varepsilon=1$) and the gain $\gamma=0.05$. It is important to note that theoretical analyses typically assume sufficiently high control frequencies. However, the robot’s physical constraints prevent operation at large frequency ranges. So, to adjust control bounds to the admissible range $|u_{1}(t)|\leq 0.22$, $|u_{2}(t)|\leq 2.84$, we select relatively small values of $\omega$ and $\gamma$. Moreover, since larger admissible values are allowed for $u_{2}$, we choose $k_{1}<k_{2}$. This choice avoids the need to further reduce $\omega$ and $\gamma$ while still satisfying the input constraints.
For the experimental validation, we use the Lyapunov function candidate (13) and controls of the form (3), (5) with one of the following sets of parameters:
$\alpha$ $=1,\;k_{1}=0.5,\;k_{2}=8$ (P1)
$\alpha$ $=1,\;k_{1}={1}/{\sqrt{2}},\;k_{2}=4\sqrt{2}$ (P2)
$\alpha$ $=4,\;k_{1}=0.5,\;k_{2}=8$ (P3)
$\alpha$ $=10,\;k_{1}=0.5,\;k_{2}=8$ (P4)
@@
Comparing the continuous-time and sampling approach, we can see that they generate clearly distinct trajectories. The continuous time approach follows a relatively direct and steep path which can especially seen from the parameter set $(P1)$, where the robot moves straight $\pi/4$ to the goal without chattering from about $1/3$ of the distance. The sampling approach produces a more curved and more chattering but faster path. When comparing the same approach across parameter sets $(P1)$ and $(P2)$, the overall trajectory shape is preserved, but parameter set $(P2)$ exhibits higher chattering for both approaches, thus $k_{1}$ and $k_{2}$ mostly influence chattering. Across the parameter sets, the two approaches maintain their characteristic differences. Both continuous and sampling approaches have a higher convergence time for parameter set $(P2)$ due to the increased chattering whereas this is more visible for the continuous time approach as the robot was almost not moving in a straight line to the target.
@@ 5 Conclusions and Future Work
The main contribution of this work is twofold. First, the possibility of practical application of the gradient-based controllers for tracking a real mobile robot toward a prescribed target position is justified. It is remarkable to note that although the theoretical stability proof is valid for sufficiently small values of the parameter $\varepsilon$ (and therefore for sufficiently large control frequency $\omega$), our experiments demonstrate the applicability of these controllers at moderate frequencies with $\varepsilon=1$. The choice of control coefficients $k_{1}$ and $k_{2}$ is justified in the present work based on control constraints for differential-drive robots combined with gradient-based control formulas under a realistic choice of a Lyapunov function candidate.
@@
A further research direction could focus on studying the well-posedness of the associated optimization problems and on developing efficient computational schemes for constructing Lyapunov function candidates. We leave this open theoretical question for future investigation.
Another promising direction concerns experimental research aimed at achieving guaranteed finite-time practical stability. In this context, by exploiting non-differentiable Lyapunov function candidates, we expect to characterize the time required to reach a prescribed neighborhood of the target point using gradient-flow-type systems. This experimental study is left for future development, possibly involving non-Lipschitz generating potential functions.
Acknowledgement
Alexander Zuyev gratefully acknowledges the funding by the European Regional Development Fund (ERDF) within the programme Research and Innovation — Grant Number ZS/2023/12/182138.

### v2

@@ 2 Stabilizing Controls for the Unicycle Kinematics
In this paper, we consider the well-known kinematic model of a unicycle rolling on the $(x_{1},x_{2})$-plane, described by the following system (see, e.g., [5]):
@@
where the vector fields $f_{1},f_{2}:{\mathbb{R}}^{3}\to{\mathbb{R}}^{3}$ are defined as $f_{1}(x)=\big(\cos x_{3},\sin x_{3},0\big)^{\top}$, $f_{2}(x)=\big(0,0,1\big)^{\top}$. Here, $(x_{1},x_{2})$ are the coordinates of the contact point of the wheel, and $x_{3}$ is the angle between the plane of the wheel and the $x_{1}$-axis. The controls $u_{1}$ and $u_{2}$ describe the translational and angular velocities of the unicycle, respectively. To simplify the presentation, all variables of this control system are assumed to be dimensionless. For the purpose of control design, we introduce the matrix ${\cal F}(x)$ composed of the vector fields $f_{1}(x)$, $f_{2}(x)$ and their Lie bracket $f_{1},f_{2}](x)=\frac{\partial f_{2}(x)}{\partial x}f_{1}(x)-\frac{\partial f_{1}(x)}{\partial x}f_{2}(x)$, and compute ${\cal F}^{-1}(x)$:
${\cal F}(x)=(f_{1},f_{2},[f_{1},f_{2}])(x)=\begin{pmatrix}\cos x_{3}&0&\sin x_{3}\\ \sin x_{3}&0&-\cos x_{3}\\ 0&1&0\end{pmatrix},{\cal F}^{-1}(x)=\begin{pmatrix}\cos x_{3}&\sin x_{3}&0\\ 0&0&1\\ \sin x_{3}&-\cos x_{3}&0\end{pmatrix}$. (2)
The nonsingularity of ${\cal F}(x)$ implies that system (1) satisfies the Lie Algebra Rank Condition (LARC): ${\rm span}(f_{1}(x),f_{2}(x),[f_{1},f_{2}](x))={\mathbb{R}}^{3}$ for all $x\in{\mathbb{R}}^{3}$. Hence, by the Chow–Rashevsky theorem, the driftless control-affine system (1) is completely controllable in ${\mathbb{R}}^{3}$.
For a given real-valued potential function $V\in C^{2}({\mathbb{R}}^{3})$, the approximate motion of system (1) in the direction of the negative gradient of $V(x)$ is generated by control functions $u^{\omega}(x,t)=(u_{1}^{\omega}(x,t),u_{2}^{\omega}(x,t))^{\top}$ of the following form [10]:
@@ 3 Justification for the Lyapunov Function Candidate
A natural way of defining an optimal Lyapunov function candidate for system (1) in $X$ can be as follows: find a function $V\in C^{2}(X;{\mathbb{R}})$ that minimizes the cost (9) under the constraint that $V$ is positive definite in $X$. This optimization problem is highly nonlinear and nonconvex, and it poses significant analytical and computational challenges. In particular, the set of positive definite functions is nonconvex in any reasonable function space, and the functional (9) exhibits singular behavior at critical points of $V$ due to the presence of $\nabla V$. Note that, in the case $U={\mathbb{R}}^{2}$, the functional $J_{X}[V]$ is invariant under multiplication of $V$ by a nonzero constant. Therefore, a solution to the above optimization problem (if it exists) is not unique without additional constraints. A comprehensive analysis of this general problem is beyond the scope of the present paper. Instead, we estimate $J_{X}[V]$ using a selected family of positive definite quadratic forms.
@@
where $\alpha$ is a parameter. In particular, computing the cost (9) for $V_{\alpha}$ with $\alpha=2,4,10$, we obtain: $J_{X}[V_{2}]\approx 0.1403,\;J_{X}[V_{4}]\approx 0.0962,\;J_{X}[V_{10}]\approx 0.0906$.
The obtained sequence of decreasing cost values justifies the use of the Lyapunov function candidates (13) for further experiments. It should be emphasized that quadratic forms cannot serve as control-Lyapunov functions for systems of the form (1) due to Brockett’s necessary stabilizability condition [6, 18]. Indeed, Brockett’s condition and Artstein’s theorem [4] imply that the considered class of nonholonomic systems does not admit any continuously differentiable control-Lyapunov function.
@@ 4.1 Experimental Setup
Robotic Platform. The experiments were conducted using a TurtleBot3 Burger (TB3) – an affordable differential-drive mobile robot (see Figure 1, left) [1]. The robot’s control and communication were handled by its onboard Raspberry Pi 4B, running Ubuntu Server 22.04.5 LTS (64-bit) and ROS 2 Humble. Mechanically, the TB3 features two off-centered, independently actuated wheels and a caster ball; thus, the robot is not a perfect unicycle. However, a differential-drive robot can be modeled as a unicycle by mapping the velocities of the left and right wheels, $v_{l}$ and $v_{r}$ (linear velocities at the wheel edge), to the unicycle velocities $u_{1}$ and $u_{2}$ introduced in (1). Let $d$ be the distance between the two wheels of the differential drive robot, and assume that the robot’s center of rotation is at the midpoint between the wheels. Then, the unicycle’s translational and angular velocities, respectively $u_{1}$ and $u_{2}$, are given by $u_{1}=\frac{v_{r}+v_{l}}{2},\quad u_{2}=\frac{v_{r}-v_{l}}{d}.$ The inverse of this transformation is handled automatically by the diff_drive_controller of ROS 2. It accepts unicycle-style Twist commands on /cmd_vel and computes the corresponding wheel velocities. The robot’s motion is constrained by its kinematics with the maximum translational velocity of 0.22 m/s and the maximum angular velocity of 2.84 rad/s. No onboard sensors were used for state estimation or control, but four reflective markers were mounted for the motion capture system.
In practice, the TurtleBot3 Burger exhibits limitations beyond velocity constraints, including motor dead zones and nonlinearities. Although no encoder or motor velocity feedback was used to compensate for these effects, the experimental results show that the proposed controller still drives the system close to the desired state in a stable manner. Incorporating additional feedback into the control law could further improve the performance in more demanding scenarios and is left for future work.
Motion Capture System. Robot pose information is obtained using an OptiTrack motion capture system, consisting of eight infrared cameras and the Motive tracking software. As previously stated, four reflective markers are attached to the top of the TB3, allowing the estimation of its $x$- and $y$-position as well as its yaw orientation. The Motive software typically computes the center point of these markers and uses it as a reference pose. However, this point does not align precisely with the robot’s pivot point. To ensure accurate control, the reference pose is adjusted manually to match the true center of rotation (see Figure 1, right). We use a dedicated wireless access point to connect to the OptiTrack system via LAN, ensuring that the TB3 and a remote laptop used for robot control can communicate over a known, low-latency wireless network.
The robot moves on a standard office carpet floor, which minimizes slip. While the performance may vary on other surfaces, the conducted experiments demonstrate the feasibility of the controller.
Another important note is that the experiments rely on the OptiTrack system for high-precision position feedback. In environments without such a system, odometry-based approaches (i.e., relying only on IMU sensors and motor encoders) would be significantly more challenging.
@@
Software Architecture. The control logic for both simulation and real-world experiments is implemented as a single ROS 2 node using the rclpy Python library. In real-world experiments, the controller subscribes to the robot pose data provided by the OptiTrack system via the NatNet protocol, whereas in simulations it subscribes to the standard /odom ROS 2 topic published by Gazebo. Control commands are published on the standard /cmd_vel topic, recognized by both the real-world TB3 and the Gazebo simulation. This allows us to use one and the same controller logic in different domains.
Although the physical dynamics of the robot are modeled in continuous time, a digital implementation of both controllers (6) and (7) cannot be entirely continuous, as delays are introduced by data retrieval, computation, and data transmission. The control commands for the TB3 are generated on a PC and sent to the robot as quickly as possible to implement strategy (7). Timestamp measurements indicate that a new control command is issued approximately every 0.5 ms, providing a high-frequency approximation of continuous-time behavior. Therefore, for notational convenience, we still refer to this implementation of controller (7) as continuous-time in the sequel.
Experimental data and controller parameters are recorded automatically for each trial. Plots were generated using Python’s matplotlib library, and all data was exported in .csv format for further analysis.
@@
Simulation Environment. Simulation experiments are conducted using the predefined TB3 Gazebo environment provided in the Robotis e-Manual [2]. An empty world is used to replicate a simple, obstacle-free arena. The same TB3 model (Burger) as in real world is used to ensure consistency in dimensions, kinematics and dynamics. The model incorporates basic rigid body physics including link masses and joint properties allowing the robot to respond realistically to velocity commands although small inaccuracies in terms of friction and inertia compared to the real-world robot may exist. Due to the software architecture the same Python controller could be used.
@@ 4.2 Experimental Results
For all experiments, the initial state is set to $x^{0}=(-0.5,-0.5,0)^{\top}$, and the target state is $x^{*}=(0,0,0)^{\top}$. The robot stops automatically if it is in the vicinity of that goal point to obtain comparable convergence time results. More precisely, an orientation angle tolerance of $\Delta x_{3}=0.05~rad$ was chosen for all parameter sets (see (P1), (P2), (P3), (P4)). While for (P1) and (P2) we chose a target distance tolerance of $\sqrt{x_{1}^{2}+x_{2}^{2}}\leq 0.005~m$, for (P3) and (P4) we chose a target distance tolerance of $\sqrt{x_{1}^{2}+x_{2}^{2}}\leq 0.03~m$ due to longer convergence times. The translational and angular velocities are governed by the control law (3), (5) with the frequency parameter $\omega=2\pi$ (i.e., $\varepsilon=1$) and the gain $\gamma=0.05$. It is important to note that theoretical analyses typically assume sufficiently high control frequencies. However, the robot’s physical constraints prevent operation at large frequency ranges. So, to adjust control bounds to the admissible range $|u_{1}(t)|\leq 0.22$, $|u_{2}(t)|\leq 2.84$, we select relatively small values of $\omega$ and $\gamma$. Moreover, since larger admissible values are allowed for $u_{2}$, we choose $k_{1}<k_{2}$. This choice avoids the need to further reduce $\omega$ and $\gamma$ while still satisfying the input constraints.
For the experimental validation, we use the Lyapunov function candidate (13) and controls of the form (3), (5) with one of the following sets of parameters: $\alpha$ $=1,\;k_{1}=0.5,\;k_{2}=8,$ (P1) $\alpha$ $=1,\;k_{1}=\tfrac{1}{\sqrt{2}},\;k_{2}=4\sqrt{2},$ (P2) $\alpha$ $=4,\;k_{1}=0.5,\;k_{2}=8,$ (P3) $\alpha$ $=10,\;k_{1}=0.5,\;k_{2}=8.$ (P4)
@@
Comparing the continuous-time and sampling approach, we can see that they generate clearly distinct trajectories. The continuous time approach follows a relatively direct and steep path which can especially be seen from the parameter set $(P1)$, where the robot moves straight $\pi/4$ to the goal without chattering from about $1/3$ of the distance. The sampling approach produces a more curved and more chattering but faster path. When comparing the same approach across parameter sets $(P1)$ and $(P2)$, the overall trajectory shape is preserved, but parameter set $(P2)$ exhibits higher chattering for both approaches, thus $k_{1}$ and $k_{2}$ mostly influence chattering. Across the parameter sets, the two approaches maintain their characteristic differences. Both continuous and sampling approaches have a higher convergence time for parameter set $(P2)$ due to the increased chattering whereas this is more visible for the continuous time approach as the robot was almost not moving in a straight line to the target.
@@ 5 Conclusions and Future Work
The main contribution of this work is twofold. First, the possibility of practical application of the gradient-based controllers for tracking a real mobile robot toward a prescribed target position is justified. It is remarkable to note that although the theoretical stability proof is valid for sufficiently small values of the parameter $\varepsilon$ (and therefore for sufficiently large control frequency $\omega$), our experiments demonstrate the applicability of these controllers at moderate frequencies with $\varepsilon=1$. The choice of control coefficients $k_{1}$ and $k_{2}$ is justified in the present work based on control constraints for differential-drive robots combined with gradient-based controls under a realistic Lyapunov function candidate.
@@
A further research direction could focus on studying the well-posedness of the associated optimization problems and on developing efficient computational schemes for constructing Lyapunov function candidates. We leave this open theoretical question for future investigation. Another promising direction concerns experimental research aimed at achieving guaranteed-time practical stabilization to reach a prescribed neighborhood of the target point, possibly exploiting non-differentiable Lyapunov function candidates.